Circular No. 8205

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ie IN NGC 4051 R. Arbour, South Wonston, U.K., reports his discovery of an apparent

supernova (mag ≈ 15.2) on an unfiltered CCD image taken with a 0.3-m f/6.3 Schmidt-Cassegrain reflector on Sept. 19.839 UT. The new object is located at α = 12^{h}03^{m}18^{s}.15, δ = +44^{o}31′34′′.6 (equinox 2000.0), which is 91′′ east and 18′′.6 south of the center of NGC 4051. SN 2003ie is not present on Digitized Sky Survey blue (1990) or red (1996) plates. T. Boles, Coddenham, England, obtained confirming unfiltered CCD images (0.35-m reflector) of the new object (at mag 14.9) on Sept. 20.173 UT, measuring a position identical with that given above.

NOVA IN M31 Boles also reports the discovery of an apparent nova (mag 18.1) in M31

on his unfiltered CCD images taken on Sept. 17.160 and 20.054 UT in the course of the U.K. Nova/Supernova Patrol. The new object is located at α = 0^{h}42^{m}59^{s}.84, δ = +41^{o}03′33′′.9 (equinox 2000.0), which is ≈ 175′′ east and 755′′ south of the center of M31. O. Trondal, Oslo, obtained a confirmation image on Sept. 20.119 with a 0.25-m reflector. The suspect is not present on Boles’ images from 2002 Sept. 28 (limiting mag 20.0) or 2003 Aug. 2 and 31 (limiting mag 19.5), and it is not present on 1989 Palomar Sky Survey red or blue plates.

SUPERNOVAE 2003hu AND 2003hz Further to IAUC 8204, T. Matheson et al. report that a spectrum of SN

2003hu (cf. IAUC 8196), obtained by M. Calkins on Sept. 19.20, shows it to be a type-Ia supernova near maximum. Narrow emission lines superposed on the spectrum of the supernova indicate a redshift of 0.075 for the host galaxy; using this value, the supernova expansion velocity is ∼ 10600 km/s for Si II (rest 635.5 nm). The spectral-feature age (Riess et al. 1997, A.J. 114, 722) of the supernova is 4 ± 2 days before maximum light. A spectrum of SN 2003hz (cf. IAUC 8201), obtained by Calkins on Sept. 19.51, shows it to be a type-Ia supernova past maximum. Narrow emission lines superposed on the spectrum of the supernova indicate a recession velocity of 6100 km/s for the host galaxy; using this value, the supernova expansion velocity is ∼ 9400 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 7 ± 2 days after maximum light.

2003 September 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green